• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Higinbotham

    Log entry time 08:59:36 on December21,2003

    Entry number 114846

    keyword=OTR and Interferometer

    Shown are the OTR and Interferometer with a 15 uA beam.
    (Note the message that the current is low is due to the
    fact EPICS is pointing to the wrong variable and is not
    seeing any current)

    The OTR is saturated (adding the filters didn't work as expected)
    but both devices are seeing the beam size
    as approx. 0.4 mm which is apporx. an energy spread of 1E-4. This
    is about what we expect with the present inject setup which has
    been optimized for G0. Tests have led accelerator to believe they
    can deliver 2.5E-5 energy spread beam to Hall A along with deliving
    G0 beam, but it has not been done for our present running conditions.

    FIGURE 1

    FIGURE 2